# Magnetic Dipole Portal Vector Dark Matter at Fixed-Targets

### v1

@@ Abstract
We present a model featuring a sub-GeV vector dark matter by augmenting the Standard Model with a new non-Abelian dark $SU(2)_{D}$, spontaneously broken by the vacuum expectation values of a scalar doublet and a triplet. Interactions between the dark and visible sectors arise through a dimension-5 non-Abelian avatar of kinetic mixing portal, inducing effective magnetic dipole couplings of the dark matter, with the photon and Z boson. The resulting spectrum of the dark gauge bosons naturally exhibits an inverse mass hierarchy between the dark matter and the $Z^{\prime}$, leading to interesting phenomenology at fixed target experiments such as LDMX through dark off-shell bremsstrahlung, dark Higgs-strahlung, invisible vector meson decay, and visible decays. We compute the thermal relic abundance across sub-GeV dark matter masses, with regions of freeze-out proceeding via forbidden annihilation into dark sector states or direct annihilation into Standard Model states. Bounds from a broad set of laboratory probes, along with cosmological and astrophysical observations, are incorporated in our analysis. Among them, the most restrictive bounds originate from direct detection experiments, Big Bang Nucleosynthesis, collider searches, and the CMB. Our results demonstrate that a sizeable region of the parameter space remains consistent with the observed relic abundance and the existing experimental results.
@@ 1 Introduction
The quest to uncover the elusive nature of dark matter (DM), strongly supported by evidence across a wide range of cosmological and astrophysical observations Balazs:2024uyj , continues along several complementary experimental and theoretical directions. Major global efforts are underway to: (i) perform direct detection searches aiming to observe interactions of the Galactic DM at the terrestrial detectors; (ii) infer DM properties through its imprint on cosmological observables such as the cosmic microwave background (CMB); (iii) identify possible astrophysical signals of DM annihilation or decay into Standard Model (SM) particles via indirect detection probes; (iv) constrain DM interactions and their effects on the SM through high-precision measurements of the SM parameters; and (v) explore the production of DM and dark-sector states in collider experiments.
@@
Along many of these fronts, the weakly interacting massive particle (WIMP) Arcadi:2017kky ; Steigman:1984ac has long served as the leading dark matter (DM) candidate, primarily due to its ability to naturally reproduce the observed relic abundance through weak-scale interactions. Although the WIMP paradigm remains appealing, growing attention has turned toward alternative possibilities. In particular, a broad class of models now considers DM candidates with masses below the GeV scale. Such light DM scenarios require the presence of a new mediator particle to facilitate interactions between the DM and the SM, thereby avoiding the Lee–Weinberg bound—an upper mass limit for thermal relics with weak-scale couplings Lee:1977ua .
@@
Since sub-GeV DM is too light to produce detectable nuclear recoils in conventional direct detection experiments, novel search strategies have been developed. A prominent approach involves direct detection of DM–electron scattering in target materials Mitridate:2022tnv ; DAMIC-M:2025luv , which provides optimal sensitivity to light DM. Concurrently, accelerator-based experiments offer complementary avenues to probe light DM, either through missing energy or momentum signatures, direct DM scattering in downstream detectors, or via visible decays of the mediator into SM particles Izaguirre:2015yja ; Krnjaic:2022ozp . Several current and upcoming experiments are specifically designed to explore this light DM regime Balan:2024cmq .
@@
Fixed-target experiments are particularly well suited for probing light dark matter (DM) owing to their well-characterized backgrounds, high luminosities, and favorable forward kinematics Berlin:2020uwy . The forthcoming Light Dark Matter eXperiment (LDMX)—a missing-momentum experiment with a planned 8 GeV electron beam incident on a tungsten target—is projected to substantially extend the current sensitivity reach LDMX:2025pkp . Theoretical studies relevant to LDMX and similar light DM searches have predominantly focused on simplified models featuring a dark photon mediator ($A^{\prime}$) Berlin:2018bsc as well as spin-1 DM scenarios encompassing both simplified and extended frameworks Catena:2023use . More recently, dark photon models incorporating higher-order dark electromagnetic moments have also been explored Catena:2025fsl .
@@
In general, any heavy new particle that can be produced in electron–nuclear collisions, such as through dark bremsstrahlung, and subsequently escape detection due to its feeble interactions with the SM, constitutes a viable target for LDMX. This broad experimental reach renders LDMX a promising platform to explore a diverse landscape of theoretical possibilities within the dark sector. Among the many possibilities, spin-1 (vector) dark matter (DM) Hambye:2008bq ; Choi:2019zeb ; Hu:2021pln ; Nomura:2020zlm ; Nomura:2021aep ; Nomura:2021tmi ; Ramos:2021txu ; Abe:2020mph ; Chu:2023zbo has been explored in diverse contexts, ranging from direct detection phenomenology Krnjaic:2023nxe ; Catena:2019hzw ; Catena:2018uae to gravitational-wave signatures Benincasa:2025tdr ; Belyaev:2025zse and freeze-in production mechanisms Krnjaic:2022wor . Previous studies, such as Ref. Catena:2023use , have focused on the on-shell production of spin-1 DM at fixed-target facilities like LDMX.
@@
Motivated by this complementary and relatively unexplored phenomenology, we propose a theoretical framework for a light spin-1 dark matter (DM) candidate. In this construction, the SM gauge group is extended by a new non-Abelian gauge symmetry, $SU(2)_{D}$ [^1] In contrast to previous studies that introduce an additional $SU(2)\times U(1)$ gauge structure Choi:2019zeb ; Ramos:2021txu , the present framework contains no extra $U(1)$ factor., which breaks spontaneously via the vacuum expectation values (vevs) of two scalar fields charged under the $SU(2)_{D}$: a triplet, $\Sigma_{D}$, and a doublet, $\Phi_{D}$. The corresponding massive gauge bosons form a triplet, comprising the DM particle $X^{\pm}_{\mu}$, and the mediator $Z^{\prime}_{\mu}$. We note that the model we present is complementary to the recent work Foguel:2025hio , where a vector DM scenario with similarities is considered.
@@
Building on the theoretical framework outlined in Section 2, we explore the phenomenology of the model in the sub-GeV mass range, considering constraints and signatures across various experimental frontiers. The DM relic abundance is computed under the assumption of thermal freeze-out, as discussed in Section 3. In Section 4, we examine existing constraints on the model parameter space arising from the collider searches, electroweak precision tests, cosmological observations, and astrophysical probes. The projected sensitivities of fixed-target experiments, with particular emphasis on LDMX LDMX:2025pkp and NA64 NA64:2025ddk , are discussed in Section 5. In Section 6, all relevant constraints are placed alongside thermal targets, revealing the complementarity of all bounds and the parameter regions yet to be excluded, followed by concluding remarks in Section 7.
@@ 2 The Model
$v_{D}\equiv=\sqrt{v_{\Phi}^{2}+2v_{\Sigma}^{2}}\,,\quad{\rm and}\quad\tan\beta\equiv\frac{v_{\Phi}}{\sqrt{2}v_{\Sigma}}\,.$ (6)
The triplet vev contributes solely to the masses of $X^{\pm}_{\mu}$, while the doublet vev generates masses for both $X^{\pm}_{\mu}$ and $X^{0}_{\mu}$. This distinction arises because $X^{\pm}_{\mu}$ carry charge under a global $U(1)_{D}$ symmetry associated with the $\tau^{3}$ generator of $SU(2)_{D}$. The doublet vev spontaneously breaks this $U(1)_{D}$, leaving behind a residual discrete $\mathbb{Z}_{2}$ symmetry that acts as a stabilizing symmetry for the dark matter. The $U(1)_{D}$ charges of the fields are summarized in the last column of Table 1. We identify the dark-charged gauge bosons $X^{\pm}_{\mu}$ as the lightest states carrying nontrivial $\mathbb{Z}_{2}$ charge, and therefore as the stable dark matter candidates in this framework. It is worth noting that this framework naturally realizes an inverse mass hierarchy between the dark matter and the neutral gauge boson $X^{0}_{\mu}$, a distinctive feature that gives rise to interesting phenomenological consequences, as elaborated in the subsequent sections.
The physical scalar spectrum involves a dark-charged scalar and two neutral scalars. Notably, we neglect the Higgs portal interaction terms in the scalar potential, since the thermal freeze-out of dark matter via Higgs portal receives stringent bounds from multiple observables, such as Higgs invisible decays, modifications to the triple-Higgs coupling, and di-Higgs production channels Biekotter:2022ckj ; ATLAS:2024ish ; CMS:2024awa .
@@
where $X_{\mu\nu}\equiv\sigma^{a}X^{a}_{\mu\nu}$, and $B_{\mu\nu}$ is the hypercharge field strength tensor and $\zeta$ parametrizes the strength of the interaction, adapting the notation from Choi:2019zeb . When $\Sigma_{D}$ receives vev, a kinetic mixing between the $X^{0}_{\mu}$ and the hypercharge gauge boson is generated as
@@
which leads to the mixing of $X^{0}_{\mu}$ with the SM neutral gauge bosons. We label the resulting BSM mass eigenstate as $Z^{\prime}$. In addition to the kinetic mixing, $\mathcal{L}_{\text{int}}$ induces contact interactions of the DM with the $Z$ boson and photon via magnetic dipole operators, as given by
@@
The parameter $\rho$ is defined as,
$\rho\equiv\frac{1}{2}\arctan\Bigg[\frac{m_{Z}^{2}\sin\theta_{w}\sin(2\zeta)}{m_{\text{DM}}^{2}\cos^{2}\beta-m_{Z}^{2}(\cos^{2}\zeta-\sin^{2}\theta_{w}\sin^{2}\zeta)}\Bigg].$ (11)
The above interactions play the key role to determine the phenomenology of the dark matter. For our analysis, we adopt a representative benchmark point with $\cos\beta=\sin\theta_{t}=0.7$, where $\theta_{t}$ denotes the mixing angle between the neutral scalar states. The masses of the dark scalars are chosen to lie close to the DM mass scale, $m_{\Sigma_{D}^{\pm}}=m_{\Sigma_{D}^{0}}=m_{\Phi_{D}^{0}}=1.2\,m_{\text{DM}}$, but remain sufficiently heavy to kinematically forbid DM decay into $\Sigma_{D}^{\pm}$.
@@ 3 Freeze-out Relic Density
The relic density of DM, measured by Planck Planck:2018vyg to be $\Omega_{\text{DM,obs}}h^{2}=0.120\pm 0.001$, is enforced in our analysis in the so-called relic targets, i.e. the contours in the parameter space that reproduce this observed value. The Boltzmann equation governing the cosmological DM number density, $n_{\text{DM}}$, is given by,
@@
where $H$ is the Hubble rate, and $\langle\sigma v_{\rm rel}\rangle$ is the thermally averaged cross section of DM annihilation. We have implemented the model in FeynRulesAlloul:2013bka , computed thermally averaged cross sections and solve the Boltzmann equation under the freeze-out approximation Gondolo:1990dk to calculate the relic density in terms of the model parameters, using an in-house developed Boltzmann solver in conjunction with micrOMEGAsBelanger:2013oya . For sub-GeV DM, annihilations into hadronic final states become important because, below the QCD confinement temperature, quarks no longer exist as free particles Izaguirre:2015zva ; Ilten:2018crw ; PDG . We therefore incorporate these hadronic channels when determining the relic targets, where they appear as resonances at DM masses satisfying $m_{\text{DM}}\simeq m_{H}/2$, with $m_{H}$ the mass of the corresponding hadron.
@@
In this model, DM may annihilate into either dark-sector or SM particles, with the dominant channels determined by the parameter regime. For small values of $\zeta$, the model enters the forbidden regime DAgnolo:2015ujb , in which DM predominantly annihilates into dark-sector states. In this case, the process shown on the left of Fig. 1 controls the relic abundance. Additional channels, such as $^{+}X^{-}\to Z^{\prime}\,\Sigma_{D}^{0}, X+X$$^{+}X^{-}\to Z^{\prime}\,\Phi_{D}^{0}$, and co-annihilations like $X^{+}\Sigma_{D}^{-}\to Z^{\prime}\Phi_{D}^{0}$ and $X^{+}\Sigma_{D}^{-}\to Z^{\prime}Z^{\prime}$, are also included, though they contribute less than 10% to the total thermally averaged cross section $\langle\sigma v_{\rm rel}\rangle$.
@@
3.1 Thermalization Condition
In the freeze-out scenario, the dark matter must remain in thermal equilibrium with the SM plasma prior to the decoupling Gondolo:1990dk . Chemical equilibrium is maintained through annihilation (production) into (from) SM species, while kinetic equilibrium is achieved by elastic scattering with SM fermions and with the $Z^{\prime}$. The corresponding interaction rate, $R\equiv n\,\langle\sigma v\rangle$, must exceed the Hubble expansion rate, $H$, at temperatures above freeze-out.
In this work, we explore values of $\sin\zeta$ spanning $10^{-10}$ to $1$. Although this range includes extremely small portal couplings, the relatively large dark gauge coupling, $g_{D}\sim 10^{-4}-10^{-2}$, ensures that the condition $R>H$ is always satisfied. The validity of the standard Boltzmann equation in Eq. (12) further requires DM to remain in kinetic equilibrium with the SM bath throughout freeze-out Binder:2017rgn ; this criterion is also fulfilled across the parameter space considered here.
DM production through freeze-in Hall:2009bx ; Cosme:2021baj , rather than freeze-out, would occur for small values of $g_{D}$ – which we leave as an interesting direction for future work.
4 Current Experimental Bounds
4.1 Colliders
Dark photons decaying into SM final states, either promptly or long lived, after being produced in proton-proton or $e^{+}e^{-}$ collisions are searched for at collider experiments. We consider the limits from the null results of the experiments LHCb LHCb:2019vmc , KLOE KLOE-2:2018kqf , BaBar BaBar:2014zli , NA48/2 NA482:2015wmo , and FASER FASER:2023tle , most relevant for the sub-GeV DM mass range.
4.2 Direct Detection
Figure 2: Direct and indirect detection detection processes. Left: DM-electron scattering that induces electronic excitations at direct detection experiments. Right: DM annihilation that inject energy into the CMB through velocity independent thermally averaged cross sections.
Direct detection experiments serve as the only way to directly measure and discover the DM present in our galaxy Schumann:2019eaa  [^2] DM collider and indirect searches act as validation of potential direct detection discoveries Catena:2024yed , and as invaluable guiding tools for DM and other new physics beyond the SM.. For WIMP scale DM, nuclear recoils are the conventional search strategy Billard:2021uyg . On the other hand, sub-GeV DM, due to its light mass, would not induce a sufficiently sized nuclear recoil to be observed. However, light DM can induce sizeable electronic excitations in the detector material from DM-electron scattering Mitridate:2022tnv ; Catena:2024rym . DM-electron scattering proceeds through a t-channel diagram via a virtual photon, as drawn in Fig. 2. DM could also scatter via a virtual $Z^{\prime}$ or $Z$-boson, however we expect these contributions to be sub-leading since the momentum transfer squared is $\ll m_{Z^{\prime}},m_{Z}$.

### v2

# Light Vector Dark Matter via a Magnetic Dipole Portal: Bridging Direct Detection and Fixed-Target Searches

@@ Abstract
We present a model featuring sub-GeV vector dark matter by augmenting the Standard Model with a new non-Abelian dark $SU(2)_{D}$, spontaneously broken by the vacuum expectation values of a scalar doublet and a triplet. Interactions between the dark and visible sectors arise through a dimension-5 non-Abelian avatar of kinetic mixing portal, inducing effective magnetic dipole couplings of the dark matter, with the photon and Z boson. The resulting spectrum of the dark gauge bosons naturally exhibits an inverse mass hierarchy between the dark matter and the $Z^{\prime}$, leading to interesting phenomenology at fixed-target experiments such as LDMX through dark off-shell bremsstrahlung, dark Higgs-strahlung, invisible vector meson decay, and visible decays. We compute the thermal relic abundance across sub-GeV dark matter masses, with regions of freeze-out proceeding via forbidden annihilation into dark sector states or direct annihilation into Standard Model states. Bounds from a broad set of laboratory probes, along with cosmological and astrophysical observations, are incorporated in our analysis. Among them, the most restrictive bounds originate from direct detection experiments, Big Bang Nucleosynthesis, collider searches, and the CMB. Our results demonstrate that a sizeable region of parameter space remains consistent with the observed relic abundance and current experimental constraints, and that fixed-target searches should be considered in tandem with direct detection and cosmological probes for an effective and comprehensive search strategy, especially in the off-shell regime.
@@ 1 Introduction
The quest to uncover the elusive nature of dark matter (DM), strongly supported by evidence across a wide range of cosmological and astrophysical observations Balazs et al. (2024), continues along several complementary experimental and theoretical directions. Major global efforts are underway to: (i) perform direct detection searches aiming to observe interactions of the Galactic DM at terrestrial detectors; (ii) infer DM properties through its imprint on cosmological observables such as the cosmic microwave background (CMB); (iii) identify possible astrophysical signals of DM annihilation or decay into Standard Model (SM) particles via indirect detection probes; (iv) constrain DM interactions and their effects on the SM through high-precision measurements of the SM parameters; and (v) explore the production of DM and dark sector states in collider experiments.
@@
Along many of these fronts, the weakly interacting massive particle (WIMP) Arcadi et al. (2018); Steigman and Turner (1985) has long served as the leading DM candidate, primarily due to its ability to naturally reproduce the observed relic abundance through weak-scale interactions. Although the WIMP paradigm remains appealing, growing attention has turned toward alternative possibilities. In particular, a broad class of models now considers DM candidates with masses below the GeV scale. Such light DM scenarios require the presence of a new mediator particle to facilitate interactions between the DM and the SM, thereby avoiding the Lee–Weinberg bound—an upper mass limit for thermal relics with weak-scale couplings Lee and Weinberg (1977).
@@
Since sub-GeV DM is too light to produce detectable nuclear recoils in conventional direct detection experiments, novel search strategies have been developed. A prominent approach involves direct detection of DM–electron scattering in target materials Mitridate et al. (2023); Aggarwal and others (2025), which provides optimal sensitivity to light DM. Concurrently, accelerator-based experiments offer complementary avenues to probe light DM, either through missing energy or momentum signatures, direct DM scattering in downstream detectors, or via visible decays of the mediator into SM particles Izaguirre et al. (2015); Krnjaic and others (2022). Several current and upcoming experiments are specifically designed to explore this light DM regime Balan and others (2025).
@@
Fixed-target experiments are particularly well suited for probing light DM owing to their well-characterized backgrounds, high luminosities, and favorable forward kinematics Berlin et al. (2020). The forthcoming Light Dark Matter eXperiment (LDMX)—a missing-momentum experiment with a planned 8 GeV electron beam incident on a tungsten target—is projected to substantially extend the current sensitivity reach Appert and others (2025). Theoretical studies relevant to LDMX and similar light DM searches have predominantly focused on simplified models featuring a dark photon mediator ($A^{\prime}$) Berlin et al. (2019) as well as spin-1 DM scenarios encompassing both simplified and extended frameworks Catena and Gray (2023). More recently, dark photon models incorporating higher-order dark electromagnetic moments have also been explored Catena et al. (2025).
@@
In general, any heavy new particle that can be produced in electron–nuclear collisions, such as through dark bremsstrahlung, and subsequently escape detection due to its feeble interactions with the SM, constitutes a viable target for LDMX. This broad experimental reach renders LDMX a promising platform to explore a diverse landscape of theoretical possibilities within the dark sector. Among the many possibilities, spin-1 (vector) DM Hambye (2009); Choi et al. (2019); Hu et al. (2021); Nomura et al. (2021); Nomura and Okada (2022, 2021); Ramos et al. (2021); Abe et al. (2020); Chu et al. (2023) has been explored in diverse contexts, ranging from direct detection phenomenology Krnjaic and Trickle (2023); Catena et al. (2019, 2018) to gravitational-wave signatures Benincasa et al. (2025); Belyaev et al. (2025) and freeze-in production mechanisms Krnjaic et al. (2023). Previous studies, such as Ref. Catena and Gray (2023), have focused on the on-shell production of spin-1 DM at fixed-target facilities like LDMX.
@@
Motivated by this complementary and relatively unexplored phenomenology, we propose a theoretical framework for a light spin-1 DM candidate. In this construction, the SM gauge group is extended by a new non-Abelian gauge symmetry, $SU(2)_{D}$ [^1] In contrast to previous studies that introduce an additional $SU(2)\times U(1)$ gauge structure Choi et al. (2019); Ramos et al. (2021), the present framework assumes only a single dark $SU(2)_{D}$ gauge symmetry, with an Abelian $U(1)_{D}$ arising solely as a residual subgroup after spontaneous symmetry breaking., which breaks spontaneously via the vacuum expectation values (vevs) of two scalar fields charged under the $SU(2)_{D}$: a triplet, $\Sigma_{D}$, and a doublet, $\Phi_{D}$. The corresponding massive gauge bosons form a triplet, comprising the DM particle $X^{\pm}_{\mu}$, and the mediator $Z^{\prime}_{\mu}$. We note that a related vector dark matter scenario has recently been explored in Foguel et al. (2025). Although both frameworks feature stable vector states arising from broken dark gauge symmetries, the underlying constructions differ, making the approaches complementary.
@@
Building on the theoretical framework outlined in Section 2, we explore the phenomenology of the model in the sub-GeV mass range, considering constraints and signatures across various experimental frontiers. The DM relic abundance is computed under the assumption of thermal freeze-out, as discussed in Section 3. The sensitivity at the direct detection experiments, viz.  DAMIC-M and PANDAX-4T, and projected sensitivities of fixed-target experiments, with particular emphasis on LDMX Appert and others (2025) and NA64 Andreev and others (2025), are discussed in Section 4 and Section 5, respectively. In Section 6, we examine existing constraints on the model parameter space arising from the collider searches, electroweak precision tests, cosmological observations, and astrophysical probes. In Section 7, all relevant constraints are placed alongside thermal targets, revealing the complementarity of all bounds and the parameter regions yet to be excluded, followed by concluding remarks in Section 8.
@@ 2 The Model
$v_{D}\equiv\sqrt{v_{\Phi}^{2}+2v_{\Sigma}^{2}}\,,\quad{\rm and}\quad\tan\beta\equiv\frac{v_{\Phi}}{\sqrt{2}v_{\Sigma}}\,.$ (6)
The triplet vev contributes solely to the masses of $X^{\pm}_{\mu}$, while the doublet vev generates masses for both $X^{\pm}_{\mu}$ and $X^{0}_{\mu}$. This distinction arises because $X^{\pm}_{\mu}$ carry charge under a global $U(1)_{D}$ symmetry associated with the $\tau^{3}$ generator of $SU(2)_{D}$. The doublet vev spontaneously breaks this $U(1)_{D}$, leaving behind a residual discrete $\mathbb{Z}_{2}$ symmetry that acts as a stabilizing symmetry for the DM. The $U(1)_{D}$ charges of the fields are summarized in the last column of Table 1. We identify the dark-charged gauge bosons $X^{\pm}_{\mu}$ as the lightest states carrying nontrivial $\mathbb{Z}_{2}$ charge, and therefore as the stable DM candidates in this framework. It is worth noting that this framework naturally realizes an inverse mass hierarchy between the DM and the neutral gauge boson $X^{0}_{\mu}$, a distinctive feature that gives rise to interesting phenomenological consequences, as elaborated in the subsequent sections.
The physical scalar spectrum involves a dark-charged scalar and two neutral scalars. Notably, we neglect the Higgs portal interaction terms in the scalar potential, since the thermal freeze-out of DM via Higgs portal receives stringent bounds from multiple observables, such as Higgs invisible decays, modifications to the triple-Higgs coupling, and di-Higgs production channels Biekötter and Pierre (2022); Aad and others (2024); Hayrapetyan and others (2025). The mass matrices and mixing angles are provided in the Appendix A.
@@
where $X_{\mu\nu}\equiv\sigma^{a}X^{a}_{\mu\nu}$, and $B_{\mu\nu}$ is the hypercharge field strength tensor and $\zeta$ parametrizes the strength of the interaction, adapting the notation from Choi et al. (2019). When $\Sigma_{D}$ receives vev, a kinetic mixing between the $X^{0}_{\mu}$ and the hypercharge gauge boson is generated as
@@
which leads to the mixing of $X^{0}_{\mu}$ with the SM neutral gauge bosons. We label the resulting BSM mass eigenstate as $Z^{\prime}$ (see Appendix A for further details). In addition to the kinetic mixing, $\mathcal{L}_{\text{int}}$ induces contact interactions of the DM with the $Z$ boson and photon via magnetic dipole operators, as given by,
@@
The parameter $\rho$ is defined in (30).
The interactions between the dark scalars, $\Sigma_{D}^{0}$ $\Sigma_{D}^{\pm}$, and $\Phi_{D}^{0}$, and the dark gauge bosons arise directly from gauge invariance and from the dimension-5 non-Abelian kinetic portal introduced in Eq. 7. The resulting feynman rules are written in Appendix C.
The above interactions play the key role to determine the phenomenology of the DM. For our analysis, we adopt a representative benchmark point with $\cos\beta=\sin\theta_{t}=0.7$, where $\theta_{t}$ denotes the mixing angle between the neutral scalar states given in Eq. 54. The masses of the dark scalars are chosen to lie close to the DM mass scale, $m_{\Sigma_{D}^{\pm}}=m_{\Sigma_{D}^{0}}=m_{\Phi_{D}^{0}}=1.2\,m_{\text{DM}}$, but remain sufficiently heavy to kinematically forbid DM decay into $\Sigma_{D}^{\pm}$.
@@ 3 Freeze-Out Relic Density
The relic density of DM, measured by Planck Aghanim and others (2020) to be $\Omega_{\text{DM,obs}}h^{2}=0.120\pm 0.001$, is enforced in our analysis in the so-called relic targets, i.e. the contours in the parameter space that reproduce this observed value. The Boltzmann equation governing the cosmological DM number density, $n_{\text{DM}}$, is given by,
@@
where $H$ is the Hubble rate, and $\langle\sigma v_{\rm rel}\rangle$ is the thermally averaged cross section of DM annihilation. We have implemented the model in FeynRulesAlloul et al. (2014), computed the thermally averaged cross sections, and solved the Boltzmann equation under the freeze-out approximation Gondolo and Gelmini (1991) to calculate the relic density in terms of the model parameters. We use an in-house developed Boltzmann solver in conjunction with micrOMEGAsBelanger et al. (2014) computed thermally averaged cross sections. For sub-GeV DM, annihilations into hadronic final states become important since below the QCD confinement temperature, quarks no longer exist as free particles Izaguirre et al. (2016); Ilten et al. (2018); et al. (Particle Data Group) (2020). We therefore incorporate these hadronic channels when determining the relic targets, where they appear as resonances at DM masses satisfying $m_{\text{DM}}\simeq m_{H}/2$, with $m_{H}$ the mass of the corresponding hadron.
@@
In this model, DM may annihilate into either dark sector or SM particles, with the dominant channels determined by the parameter regime. For small values of $\zeta$, the model enters the forbidden regime D’Agnolo and Ruderman (2015), in which DM predominantly annihilates into dark sector states. In this case, the process shown on the left of Fig. 1 controls the relic abundance. Additional channels, such as $^{+}X^{-}\to Z^{\prime}\,\Sigma_{D}^{0}, X+X$$^{+}X^{-}\to Z^{\prime}\,\Phi_{D}^{0}$, and co-annihilations like $X^{+}\Sigma_{D}^{-}\to Z^{\prime}\Phi_{D}^{0}$ and $X^{+}\Sigma_{D}^{-}\to Z^{\prime}Z^{\prime}$, are also included, though they contribute less than 10% to the total thermally averaged cross section $\langle\sigma v_{\rm rel}\rangle$.
@@
In the freeze-out scenario, the DM must remain in thermal equilibrium with the SM plasma prior to the decoupling Gondolo and Gelmini (1991). Chemical equilibrium is maintained through annihilation (production) into (from) SM species, while kinetic equilibrium is achieved by elastic scattering with SM fermions and with the $Z^{\prime}$. The corresponding interaction rate, $R\equiv n\,\langle\sigma v\rangle$, must exceed the Hubble expansion rate, $H$, at temperatures above freeze-out.
In this work, we explore values of $\sin\zeta$ spanning $10^{-10}$ to $1$. Although this range includes extremely small portal couplings, the relatively large dark gauge coupling, $g_{D}\sim 10^{-4}-10^{-2}$, ensures that the condition $R>H$ is always satisfied. The validity of the standard Boltzmann equation in Eq. (11) further requires DM to remain in kinetic equilibrium with the SM bath throughout freeze-out Binder et al. (2017); this criterion is also fulfilled across the parameter space considered here.
DM production through freeze-in Hall et al. (2010); Cosme et al. (2021), rather than freeze-out, would occur for small values of $g_{D}$ – which we leave as an interesting direction for future work.
4 Direct Detection at DAMIC-M and PANDAX-4T
Figure 2: Direct and indirect detection processes. Left: DM-electron scattering that induces electronic excitations at direct detection experiments. Right: DM annihilation that inject energy into the CMB through velocity independent thermally averaged cross sections.
Direct detection experiments serve as the only way to directly measure and discover the DM present in our galaxy Schumann (2019) [^2] DM collider and indirect searches act as validation of potential direct detection discoveries Catena et al. (2024), and as invaluable guiding tools for DM and other new physics beyond the SM.. For WIMP scale DM, nuclear recoils are the conventional search strategy Billard and others (2022). On the other hand, sub-GeV DM, due to its light mass, would not induce a sufficiently sized nuclear recoil to be observed. However, light DM can induce sizeable electronic excitations in the detector material from DM-electron scattering Mitridate et al. (2023); Catena and Spaldin (2024). DM-electron scattering proceeds through a t-channel diagram via a virtual photon, as drawn in Fig. 2. DM could also scatter via a virtual $Z^{\prime}$ or $Z$-boson, however we expect these contributions to be sub-leading since the momentum transfer squared is $\ll m_{Z^{\prime}},m_{Z}$.